OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are (10, 6) and (-4, 8). and Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1)

hartnn (hartnn):

for first, The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

so like x= 10+-4/2

hartnn (hartnn):

for 2nd question, first find the slope of line containing points P and Q, using the formula i gave you. then i'll tell u what to do next... yup, x = (10-4)/2 =... ?

OpenStudy (anonymous):

so for the first one is it (3,7)

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

yay :) for the second im still lost

hartnn (hartnn):

first find the slope of line with points P(0, -4), Q(5, 1) same formula we used in previous question, should i mention it again ?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

like -4-1/0-5

OpenStudy (anonymous):

-5/-5

hartnn (hartnn):

yup, so the slope m=1, right ?

OpenStudy (anonymous):

i tought it was -1?

hartnn (hartnn):

- *minus* in numerator and - in denominator gets cancelled, -1/-1 = 1

OpenStudy (anonymous):

oh okay i see

hartnn (hartnn):

now you have the slope(m=1) and take one point say (0,-4) --> x1=0, y1 =-4 plug these in y-y1 = m (x-x1)

OpenStudy (anonymous):

y-(-4)=1(x-0)

hartnn (hartnn):

yes, simplify ..

OpenStudy (anonymous):

so y-4=1x-0 ??

hartnn (hartnn):

the standard form of line equation is : ax+by = c (or ax+by+c=0) so, to get y-4 =x in that form, i would suggest you to first subtract x from both sides, what u get ?

OpenStudy (anonymous):

??uhm so x-y=-4

hartnn (hartnn):

yes, that would be correct :)

OpenStudy (anonymous):

yess thanks

OpenStudy (anonymous):

can you help me on this on plz Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

hartnn (hartnn):

sorry i lost connection and y-(-4) = y+4=x will give u x-y=+4